Online Store Monthly E- Commerce sales (in 1000s)Online Advertising (1000s)13681.7023401.5036652.8049545.0053311.3065562.2073761.30




1
Expert's answer
2021-11-09T11:37:53-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
Online & Monthly\ E- & Online\ Advertising \\ 
 Store & commerse\ Sales & Dollars \\
 & \ (in\ 1000s) & (1000s)\\
 \hline
 1 & 368 & 1.7\\
 2 & 340 & 1.5\\
 3 & 665 & 2.8\\
 4 & 954 & 5\\
 5 & 331 & 1.3\\
 6 & 556 & 2.2\\
 7 & 376 & 1.3\\
\end{array}"

a) Online Advertising Dollars is independent variable "(X)."

Monthly E-commerse Sales is dependent variable "(Y)."


"\bar{X}=\dfrac{1}{n}\sum_iX_i=\dfrac{15.8}{7}=2.257143"

"\bar{Y}=\dfrac{1}{n}\sum_iY_i=\dfrac{3590}{7}=512.857143"

"SS_{XX}=\sum_i(X_i-\bar{X})^2=10.537143"

"SS_{YY}=\sum_i(Y_i-\bar{Y})^2=322280.857143"

"SS_{XY}=\sum_i(X_i-\bar{X})(Y_i-\bar{Y})=1806.757143"

"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{1806.757143}{10.537143}=171.465564"

"n=\bar{Y}-m\bar{X}"

"=512.857143-171.465564(2.257143)"

"=125.834870"

Therefore, we find that the regression equation is:


"Y=125.834870+171.465564X"



b) Correlation coefficient:


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}"

"=\dfrac{1806.757143}{\sqrt{10.537143}\sqrt{322280.857143}}"

"=0.980440"

We have strong positive correlation.